At temperature of 20 degree Celsius, the length of a copper rod is 30 cm with a mass of 250g .Ravi heats up the copper rod by supplying 5 kJ of heat .Determine the length of the rod(in cm) after being heated.

1
Expert's answer
2022-06-13T08:23:51-0400
"Q=cm\Delta t"

"\Delta t=\frac{Q}{cm}=\frac{5000\:\rm J}{385\:\rm J/(K\cdot kg)*0.25\:kg}=52^\circ\rm C"

"l=l_0(1+\alpha\Delta t)\\
=30\:\rm cm(1+17*10^{-6}*52)=30.027\:cm"